OpenStudy (anonymous):

I am working on finding slopes, and I am completely confused. I have to find the slope, if there is one, of the line containing two pair of points. The points are (-1, -8) and (-6 and -16) I know I need to do rise over run which consists on y2-y1 over x2-x1. I did the work but came up with -8/-5 and I am not allowed to graph, so would that mean that the slope is undefined? Please if anyone could help I would really appreciate it.

OpenStudy (mathstudent55):

(y2 - y1) / (x2 - x1) = ( -16 - (-8) ) / ( -6 - (-1) ) = (-16 + 8) / (-6 + 1) = (-8)/(-5) = 8/5 Slope is 8/5

OpenStudy (mathstudent55):

You did your calculations correctly. All you need to do is simplify -8/-5 by dividing the numerator and denominator by -1, to get 8/5.

OpenStudy (anonymous):

I know where your answer is coming from, but why do you divide by -1?